OpenStudy (anonymous):

Is the function given by y=-x^4+2x^2-1 odd even or neither, dumb it down for me as how you find the answer please

OpenStudy (anonymous):

A function f(x) is even if \[f(-x)=f(x)\] And it is odd if \[f(-x)=-f(x)\] Now, substitute x in your function with -x and see how it works out: \[f(x)=-x^4+2x^2-1\] then \[f(-x)=-(-x)^4+2(-x)^2-1\] You know that -x = -1*x, and that -1 raised to an EVEN power will always be positive 1. So, \[f(-x)=-(-1*x)^4+2(-1*x)^2-1\] \[f(-x)=-1*(x)^4+2*1(x)^2-1\] \[f(-x)=-x^4+2x^2-1\] Therefore \[f(-x)=f(x)\] Which means that the function is even.

OpenStudy (anonymous):

Thank you, really helpful exactly what I needed